OpenStudy (anonymous):

how do you solve and graph this compound inequality : 5 less than or equal to y+ 2 less than or equal to 11

OpenStudy (queelius):

\[5 \le y+2 \le 11\] We are interested in what values y can take, so we should first solve for y. By subtracting two from each component, we get: \[5-2 \le y + 2 - 2 \le 11 - 2 = 3 \le y \le 9\] So, y is a value strictly between 3 and 9. How might you graph this?